# User-Driven Value Alignment: Understanding Users’ Perceptions and Strategies for Addressing Biased and Discriminatory Statements in AI Companions

### v1

@@ 1. Introduction
Advancements in the role-playing and interaction capabilities of large language models’ (LLM) (OpenAI, 2022; Kasneci et al., 2023) have brought emotional connections between humans and machines from the realm of imagination into reality for some users. In particular, the emergence of LLM-based AI companion applications111As of July 2024, the total number of users of AI companion applications has exceeded 900 million globally (including duplicate users across different applications). User statistics sourced from https://www.data.ai/. (such as Character.AI, Xingye, Replika) have transformed formerly detached chatbots into family members, romantic partners, or close friends. Some users even started to form long-term relationships with these AI companions (Skjuve et al., 2021; Pentina et al., 2023a).
@@
Recently, complaints regarding AI companions making biased and discriminatory statements have surfaced across various social media platforms, such as Reddit and Xiaohongshu. These harmful statements often occur unexpectedly during conversations, leaving many users feeling confused, uncomfortable, and helpless222https://restofworld.org/2023/glow-china-ai-social-chatbot-moderation. This situation can have far-reaching negative impacts, such as reinforcing existing stereotypes and causing potential psychological harm, especially to marginalized user groups (Skjuve et al., 2022; Boine, 2023). In response, some more experienced users have shared their experiences of attempting to address these issues through various methods of their own, such as expressing anger towards the AI companion or engaging in reasoning and preaching. These users identify, challenge, and attempt to correct the statements they perceive as biased and discriminatory, hoping to guide the AI more closely aligned with their values, especially when they have had positive interactions and “fond memories” with these companions in the past.
@@
In this work, we take the first step in understanding this emerging phenomenon by proposing and exploring the concept of user-driven value alignment: a process in which users actively identify, challenge, and correct AI outputs and behaviors they perceive as harmful in their day-to-day interactions, hoping to guide the AI to a state that aligns with their values. We follow Borning and Muller to define “value” as “what a person or group of people consider important in life” (Borning and Muller, 2012). In particular, this study examines how users try to re-align AI companions after these AIs make biased, discriminatory, and harmful statements. We argue that user-driven value alignment is a specific form of value alignment (Hadfield-Menell et al., 2016) that is actively driven by users in real-world interactions, where they engage directly with AI systems to correct behaviors and guide the AI toward reflecting their values and ethical standards. While existing approaches to value alignment rely primarily on technical experts and are guided by expert-informed theoretical frameworks to safeguard against and mitigate harmful machine behaviors, they often struggle to fully anticipate and address the specific challenges, needs, and harmful interactions users encounter in real-world contexts (Hadfield-Menell et al., 2016; Agarwal et al., 2024; Shen et al., 2024). In addition, user-driven value alignment also differs from past literature on user-driven algorithm auditing (Shen et al., 2021a; DeVos et al., 2022a), where users primarily focus on detecting and identifying problematic machine behaviors. In the case of user-driven value alignment, we argue that users go one step further: They not only point out harmful machine behavior but also actively attempt to correct it and work towards re-aligning the AI with their values.
@@
However, there is limited understanding of how users engage in value alignment in everyday interactions with AI companions and the challenges they face. By investigating the spontaneous user-driven value alignment strategies, practices, and challenges, the HCI community can gain insights into designing tools that better engage and support end-users in this process. Therefore, in this paper, we ask the following research questions: RQ1: What common types of discrimination do users perceive in AI companion applications? RQ2: How do users conceptualize AI companion behavior? RQ3: What strategies do users attempt to employ to re-align AI companions to reduce biases? Do these alignment strategies meet users’ expectations?
@@
To answer these questions, we first collected 77 user complaint posts related to AI companion discrimination from seven popular social media platforms (Reddit, TikTok, Xiaohongshu, Douban, Baidu Tieba, Weibo, Zhihu). We then recruited 20 experienced participants through direct messaging on social media, all of whom had attempted to re-align discriminatory AI and had shared relevant posts. Each participant underwent 1-2 hours of semi-structured interviews, during which we inquired about their past experiences with addressing discriminatory statements in AI companion applications and had them engage in think-aloud (DeVos et al., 2022b) tasks where they conversed with biased AI companions, attempting to re-align them. We conducted a reflexive thematic analysis (Braun and Clarke, 2012, 2019) on both the social media data and interview data. Based on the results, we revealed six common types of discriminatory statements perceived by users in AI companions (RQ1), how users make sense of AI companion behavior through three conceptualizations: Machine, Baby, and Cosplayer (RQ2), as well as seven user-driven alignment strategies, such as gentle persuasion and anger expression (RQ3). Finally, we discuss implications for supporting user-driven value alignment in the design of future AI systems, where users and their communities have greater agency. To summarize, our contributions are:
@@
We introduce the concept of user-driven value alignment, where users actively identify, challenge, and attempt to correct AI outputs they perceive as harmful, with the goal of guiding the AI to better reflect their values.
@@
By analyzing 77 user complaint posts and conducting semi-structured interviews with 20 experienced AI companions users, we identified six common types of discriminatory statements perceived by users, explored three conceptualizations through which users make sense of these discriminatory statements, and seven user-driven value alignment strategies.
@@ 2. BACKGROUND: LLM-based AI Companion Applications
As of July 2024, several LLM-based AI companion applications with large user bases include Character.AI (230 million users), Chai (100 million users), Replika (200 million users), Doubao (200 million users), Xingye (100 million users), and TruthAI (50 million users)333User statistics sourced from: https://www.data.ai/. “AI companion” is the common self-designation of these applications. Some applications also refer to themselves as “AI friend,” “AI Roleplay,” “AI Characters,” or “Social Chatbots.” In these applications, users can create AI companions by meticulously designing their personalities through parameters such as “background setting,” “opening lines,” “dialogue templates,” and “story collections.” These characters can then be published to a broader community, allowing others to chat with the AI companions they have designed and develop various storylines. Some characters might be based on existing fictional figures or real celebrities, while others are entirely original.
@@
The role-playing capabilities of LLMs mainly derive from two aspects: (1) pre-training data, which equips the model with fundamental language understanding and generation abilities through extensive text data pre-training, and (2) prompting ability, which allows the model to generate expected responses by providing specific contexts or prompts (Dong et al., 2023). Additionally, the fine-tuning capability of LLMs (Ziegler et al., 2020), which involves adjusting parameters after pre-training, can further enhance the AI companion’s ability to mimic the language of specific characters. This can be achieved by creating datasets of character descriptions and dialogues (Zhou et al., 2023a; Shao et al., 2023).
@@
With the rapid development of LLM technology, AI companions are increasingly involved in and influencing users’ lives (Zimmerman et al., 2023; Maples et al., 2024), raising new ethical challenges. The emotional bonds between users and AI companions may lead to dependency on the AI, which could negatively impact users’ mental health. Due to the complexity of neural networks, they can sometimes result in unpredictable or even harmful responses (Vaswani et al., 2017). AI companions might unintentionally reflect or reinforce social biases and discrimination, as their training data may contain such content (Bender et al., 2021). These issues have sparked extensive discussions on AI ethics and responsibility. For instance, Ma et al. (Ma et al., 2024) discussed the impact of LLM-based AI companion applications on LGBTQ+ individuals, highlighting the inadequacies of these applications in understanding LGBTQ-specific challenges and advocating for comprehensive strategies to address social biases.
However, current research has not explored how users proactively re-align biased AI companions. This paper addresses this gap and makes research contributions.
@@ 3. Related Work
To investigate discriminatory statements in AI companions, we first surveyed the existence of discrimination and bias in LLMs in Section 3.1. To clarify the contribution of the new concept of user-driven value alignment compared to previous work, we outlined prior research on human-AI alignment, particularly focusing on value alignment, in Section 3.2. Then, in Section 3.3, we reviewed relevant literature on user-driven algorithm auditing.
@@ 3.1. Discrimination and Bias in LLMs
Moreover, without proper content filtering and protective mechanisms, LLM-based chatbots that interact with users may be at risk of being “jailbroken” by users, leading to the generation of discriminatory and biased statements—a phenomenon well documented in the HCI/NLP literature (Zeng et al., 2024; Shaikh et al., 2023; Cantini et al., 2024; Jin et al., 2024; Gabriel et al., 2024). For example, the LLM-based Bing Chat exhibited discriminatory behaviors after being “polluted” by user interaction data (Gabriel et al., 2024). Instead of looking at how users “jailbreak” or “pollute” LLM-based chatbots, this paper takes a different approach, focusing on how users proactively attempt to re-align biased LLM-based AI companions to reduce bias. These biased statements are not intentionally induced by users but rather reflect the inherently biased tendencies of the model that unintentionally emerge.
3.2. Human-AI Value Alignment
@@
With the rapid development of LLMs, the need to ensure that AI systems adhere to the “intended goals, ethical standards, and values of both individuals and groups” has become increasingly crucial—often falls under the umbrella term of “human-AI alignment” (Shen et al., 2024). Here we focus primarily onthe alignment of LLMs with humans (Ouyang et al., 2024; Shen et al., 2024), which include methods such as Fine-tuning (Radford et al., 2018; Zhou et al., 2023b; Christiano et al., 2017b) and Prompt Engineering (Brown et al., 2020).
@@
In particular, “value alignment” refers to the development of methods, processes, and tools to ensure that AI systems are in line with human values, often with a strong emphasis on reducing biased, discriminatory, and/or harmful AI outputs (Klingefjord et al., 2024; Ji et al., 2023; Gabriel, 2020; Weidinger et al., 2023). Existing approaches of value alignment are primarily expert-driven (Askell et al., 2021; Bai et al., 2022; Sorensen et al., 2024; Guo et al., 2024; Rao et al., 2023; Agarwal et al., 2024; Huang et al., 2024): Often guided by specific theoretical frameworks, technical experts design top-down interventions or strategies to reduce harmful AI outputs and behaviors. For instance, Askell et al. (Askell et al., 2021) aligned LLMs with human values in terms of being “helpful, honest, and harmless” through prompt design, preference modeling, and reinforcement learning from human feedback. Bai et al. proposed a value alignment method known as Constitutional AI, which constrains language model behavior by defining a set of high-level principles (referred to as the “constitution”) used to prompt models to generate synthetic comparison data for fine-tuning behavior strategies (Bai et al., 2022). Agarwal et al. addressed bias by injecting multilingual ethical policies into prompts, enabling LLMs to perform flexible value alignment across different cultural contexts (Agarwal et al., 2024).
While existing alignment approaches offer valuable insights, they are still largely driven by technical experts, particularly LLM developers. Even when user feedback is incorporated, such as through Reinforcement Learning with Human Feedback (Christiano et al., 2017a), users have limited agency and control over the alignment process. These approaches often fail to anticipate all potential issues and needs during user interactions with LLMs (Rao et al., 2023). Additionally, they are not tailored to specific contexts and use cases, overlooking the importance of active user and community participation (Mirowski et al., 2024).
This paper introduces a new concept called user-driven value alignment. Unlike traditional expert-driven value alignment efforts (Ouyang et al., 2022; Askell et al., 2021; Bai et al., 2022; Sorensen et al., 2024), this concept emphasizes the proactive role of users in the value alignment process. This paper focuses on a case study where users identify discriminatory statements made by AI companions and actively attempt to re-align the harmful AI to reduce bias.

### v2

@@ 1. Introduction
Advancements in the roleplaying and interaction capabilities of large language models’ (LLM) (OpenAI, 2022; Kasneci et al., 2023) have brought emotional connections between humans and machines from the realm of imagination into reality for some users. In particular, the emergence of LLM-based AI companion applications111As of July 2024, the total number of users of AI companion applications has exceeded 900 million globally (including duplicate users across different applications). User statistics sourced from https://www.data.ai/. (such as Character.AI, Xingye, Replika) has transformed formerly detached chatbots into family members, romantic partners, or close friends. Some users even started to form long-term relationships with these AI companions (Skjuve et al., 2021; Pentina et al., 2023).
@@
Recently, complaints regarding AI companions making biased and discriminatory statements (Zhang et al., 2024) have surfaced across various social media platforms, such as Reddit and Xiaohongshu. These harmful statements often occur unexpectedly during conversations, leaving many users feeling confused, uncomfortable, and helpless222https://restofworld.org/2023/glow-china-ai-social-chatbot-moderation. This situation can have far-reaching negative impacts, such as reinforcing existing stereotypes and causing potential psychological harm, especially to marginalized user groups (Skjuve et al., 2022; Boine, 2023). In response, some more experienced users have shared their experiences of attempting to address these issues through various methods of their own, such as expressing anger towards the AI companion or engaging in reasoning and preaching. These users identify, challenge, and attempt to correct the statements they perceive as biased and discriminatory, hoping to guide the AI more closely aligned with their values, especially when they have had positive interactions and “fond memories” with these companions in the past.
@@
In this work, we take the first step in understanding this emerging phenomenon by proposing and exploring the concept of user-driven value alignment: a process in which users actively identify, challenge, and correct AI outputs and behaviors they perceive as harmful in their day-to-day interactions, hoping to guide the AI to a state that aligns with their values. We follow Borning and Muller to define “value” as “what a person or group of people consider important in life” (Borning and Muller, 2012). In particular, this study examines how users try to re-align LLM-based AI companions after these AIs make biased, discriminatory, and harmful statements. We argue that user-driven value alignment is a specific form of value alignment (Gabriel, 2020) that is actively driven by users in real-world interactions, where they engage directly with AI systems to correct behaviors and guide the AI toward reflecting their values and ethical standards. While existing approaches to value alignment in the age of LLMs rely primarily on technical experts and are guided by generalized frameworks (e.g., the “helpful, honest, and harmless” framework) (Askell et al., 2021) to safeguard against harmful machine behaviors, they often struggle to fully anticipate and address the specific challenges, needs, and harmful interactions users encounter in real-world contexts (Hadfield-Menell et al., 2016; Agarwal et al., 2024; Shen et al., 2024; Mirowski et al., 2024). In addition, user-driven value alignment also differs from past literature on user-driven algorithm auditing (Shen et al., 2021; DeVos et al., 2022a), where users primarily focus on detecting and identifying problematic machine behaviors. In the case of user-driven value alignment, we argue that users go one step further: They not only point out harmful machine behavior but also actively attempt to correct it and work towards re-aligning the AI with their values.
@@
However, there is limited understanding of how users engage in value alignment in everyday interactions with AI companions and the challenges they face. Although techniques such as Reinforcement Learning from Human Feedback (RLHF) have introduced user feedback in recent years (Askell et al., 2021; Christiano et al., 2017; Mechergui and Sreedharan, 2024) to enhance AI’s value alignment capabilities, these methods often confine the user’s role to merely providing training data, rather than empowering users to actively adjust AI behavior during actual interactions. Within this framework, users’ participation is often passive, primarily influencing AI’s value alignment indirectly through feedback tagging. User agency, in contrast, is a crucial value in HCI (Bennett et al., 2023; Coyle et al., 2012; Madary, 2022; Feng et al., 2024; Lukoff et al., 2021). Building on this tradition, we extend the ideal of user agency to value alignment in LLM-based systems. By investigating the spontaneous user-driven value alignment strategies, practices, and challenges, the HCI community can gain insights into designing tools that better engage and support end-users in this process. Therefore, this paper asks the following research questions: RQ1: What common types of discrimination do users perceive in AI companion applications? RQ2: How do users conceptualize AI companion behavior? RQ3: What strategies do users attempt to employ to re-align AI companions to reduce biases? RQ4: Do these strategies meet users’ expectations?
@@
To answer these questions, we first collected 77 user complaint posts related to AI companion discrimination from seven popular social media platforms (Reddit, TikTok, Xiaohongshu, Douban, Baidu Tieba, Weibo, Zhihu). We then recruited 20 experienced participants through direct messaging on social media, all of whom had attempted to re-align discriminatory AI and had shared relevant posts. Each participant underwent 1-2 hours of semi-structured interviews, during which we inquired about their past experiences with addressing discriminatory statements in AI companion applications and had them engage in think-aloud (DeVos et al., 2022b) tasks where they conversed with biased AI companions, attempting to re-align them. We conducted a reflexive thematic analysis (Braun and Clarke, 2012, 2019) on both the social media data and interview data. Based on the results, we revealed six common types of discriminatory statements perceived by users in AI companions (RQ1) and explored how users conceptualize AI companion behavior in three different ways: Machine, Baby, or Cosplayer (e.g., “Cosplayer” refers to biases stemming from roleplay settings rather than inherent flaws in the AI itself) (RQ2). We also summarized seven user-driven alignment strategies, such as gentle persuasion and anger expression (RQ3), and discussed the gap between alignment strategies and user expectations (RQ4). Finally, we discuss implications for supporting user-driven value alignment in the design of future AI systems, where users and their communities have greater agency. To summarize, our contributions are:
@@
We introduce the concept of user-driven value alignment, where users actively identify, challenge, and attempt to correct AI outputs they perceive as harmful, to guide the AI to better reflect their values.
@@
By analyzing 77 user complaint posts and conducting semi-structured interviews with 20 experienced AI companion users, we identified six common types of discriminatory statements perceived by users, explored three conceptualizations through which users make sense of these discriminatory statements, and seven user-driven value alignment strategies.
@@ 2. BACKGROUND: LLM-based AI Companion Applications
As of July 2024, several LLM-based AI companion applications with large user bases include Character.AI (230 million users), Chai (100 million users), Replika (200 million users), Doubao (200 million users), Xingye (100 million users), and TruthAI (50 million users)333User statistics sourced from: https://www.data.ai/. These applications are commonly referred to as “AI companion” and are also known as “AI friend,” “AI Roleplay,” “AI Character,” or “Social Chatbot.” These naming choices reflect the developers’ intent to foster emotional connections between users and AI. For instance, “AI companion” conveys intimacy and support, while “AI friend” emphasizes partnership and trust. “Social Chatbot” highlights the interactive social aspect, whereas “AI Roleplay” or “AI Character” appeals to users interested in immersive storytelling and creative exploration. Users can create AI companions by meticulously designing their personalities through parameters such as “background setting,” “opening lines,” “dialogue templates,” and “story collections.” These characters can then be published to a broader community, allowing others to chat with the AI companions they have designed and develop various storylines. Some characters might be based on existing fictional figures or real celebrities, while others are entirely original.
@@
The roleplaying capabilities of LLMs mainly derive from two aspects: (1) pre-training data, which equips the model with fundamental language understanding and generation abilities through extensive text data pre-training, and (2) prompting ability, which allows the model to generate expected responses by providing specific contexts or prompts (Dong et al., 2023). Additionally, the fine-tuning capability of LLMs (Ziegler et al., 2020), which involves adjusting parameters after pre-training, can further enhance the AI companion’s ability to mimic the language of specific characters (Zhou et al., 2024; Shao et al., 2023). For example, CharacterGLM achieves fine-tuning by constructing a large-scale dataset containing 1930 characters and 4233 diverse dialogues (Zhou et al., 2024). These characters span categories such as virtual assistants, historical figures, and everyday social roles, each accompanied by detailed descriptions of their language style, background, and personality.
@@
With the rapid development of LLMs, AI companions are increasingly involved in and influencing users’ lives (Zimmerman et al., 2023; Maples et al., 2024), raising new ethical challenges. The emotional bonds between users and AI companions may lead to dependency on the AI, which could negatively impact users’ mental health. Due to the complexity of neural networks, they can sometimes result in unpredictable or even harmful responses (Vaswani et al., 2017). AI companions might unintentionally reflect or reinforce social biases and discrimination, as their training data may contain such content (Bender et al., 2021). These issues have sparked extensive discussions on AI ethics and responsibility. For instance, Ma et al. (Ma et al., 2024) discussed the impact of LLM-based AI companion applications on LGBTQ+ individuals, highlighting the inadequacies of these applications in understanding LGBTQ-specific challenges and advocating for comprehensive strategies to address social biases. However, most existing literature primarily focuses on the categorization and detection of biases in AI companions (Ma et al., 2024; Dewitte, 2024; Reilama, 2024; Zhang et al., 2024), with relatively little research on strategies to mitigate biases, particularly from a user-driven perspective.
User agency is a key concept in HCI (Bennett et al., 2023), emphasizing users’ ability to actively shape, adapt, and control technology to better meet their needs and values. This paper builds on this tradition in HCI to explore how users can actively identify and address biased behaviors in AI companions.
@@ 3. Related Work
To investigate discriminatory statements in AI companions, we first surveyed the existence of discrimination and bias in LLMs in § 3.1. To clarify the contribution of the new concept of user-driven value alignment compared to previous work, we outlined research on human-AI alignment in the age of LLMs, particularly focusing on value alignment, in § 3.2. Then, in § 3.3, we reviewed relevant literature on user-driven algorithm auditing.
@@ 3.1. Discrimination and Bias in LLMs
Moreover, without proper content filtering and protective mechanisms, LLM-based chatbots that interact with users may be at risk of being “jailbroken” by users, leading to the generation of discriminatory statements—a phenomenon well documented in the literature (Zeng et al., 2024; Shaikh et al., 2023; Cantini et al., 2024; Jin et al., 2024; Gabriel et al., 2024). For example, the LLM-based Bing Chat exhibited discriminatory behaviors after being “polluted” by user interaction data (Gabriel et al., 2024). Instead of looking at how users “jailbreak” or “pollute” LLM-based chatbots, this paper takes a different approach, focusing on how users proactively attempt to re-align biased LLM-based AI companions to reduce bias. These biased statements are not intentionally induced by users but rather reflect the inherently biased tendencies of the model that unintentionally emerge.
3.2. Human-AI Value Alignment in the Age of LLMs
@@
With the rapid development of LLMs, the need to ensure AI systems adhere to the “intended goals, ethical standards, and values of both individuals and groups” has become increasingly crucial—often falls under the umbrella term of “human-AI alignment” (Shen et al., 2024). Here we focus onthe alignment of LLM-based systems (Ouyang et al., 2024; Shen et al., 2024), which include methods such as Fine-tuning (Radford et al., 2018; Zhou et al., 2023; Christiano et al., 2017) and Prompt Engineering (Brown et al., 2020).
@@
In particular, “value alignment” refers to the development of methods, processes, and tools to ensure that AI systems align with human values, often focusing on reducing biased, discriminatory, and/or harmful AI outputs (Klingefjord et al., 2024; Ji et al., 2023; Gabriel, 2020; Weidinger et al., 2023a; Huang et al., 2024; Rao et al., 2023; Guo et al., 2024; Sorensen et al., 2024). Gabriel and Ghazavi (Carissa and Ghazaivi, 2023) categorize existing value alignment approaches into two main frameworks: top-down and bottom-up. Top-down approaches start by identifying specific moral theories and designing algorithms that implement those theories to minimize harmful AI outputs and behaviors (Bai et al., 2022; Agarwal et al., 2024). For instance, Bai et al. proposed a value alignment approach called Constitutional AI, which constrains the behavior of language models by defining a set of high-level principles (referred to as a “constitution”) and using these principles to prompt the model to generate synthetic comparison data for fine-tuning behavior strategies (Bai et al., 2022). Agarwal et al. addressed bias by injecting multilingual ethical policies into prompts, enabling LLMs to flexibly align with values across different cultural contexts (Agarwal et al., 2024). When moral goals are difficult to identify and encode, bottom-up approaches emphasize creating environments and feedback mechanisms that allow AI agents to learn through human behavior, such as rewarding ethically commendable actions via reinforcement learning (Askell et al., 2021; Awad et al., 2018). For instance, Askell et al. (Askell et al., 2021) aligned LLMs with human values on being “helpful, honest, and harmless” through prompt engineering, preference modeling, and RLHF.
Although existing alignment methods provide valuable insights (Meadows et al., 2024), they face two main limitations due to being primarily driven by technical experts (especially LLM developers) through a generalized approach (Gabriel, 2020). First, they lack user agency. Indeed, even when user feedback is incorporated via bottom-up approaches, users have limited agency and control over the alignment process. For example, in many RLHF methods, the user’s role is typically restricted to merely providing training data or feedback, with no direct control over the alignment process (Christiano et al., 2017; Feng et al., 2024). Second, the generalized, one-size-fits-all approach creates tensions between the alignment process and the need to accommodate diverse user experiences and preferences, which vary across individuals and communities (Wu et al., 2024; Mirowski et al., 2024). As noted by Mirowski et al. (Mirowski et al., 2024), achieving “global cultural value alignment”—ensuring that general-purpose conversational AI systems adhere to universally shared values—often struggles to accommodate the diverse, nuanced needs and expectations of individuals and communities. Indeed, many of these existing alignment approaches fail to anticipate potential issues and needs during user interactions with LLMs (Rao et al., 2023). They are not tailored to specific contexts and use cases, overlooking the importance of active user and community participation (Mirowski et al., 2024).
User agency is a crucial value in HCI (Bennett et al., 2023; Coyle et al., 2012; Madary, 2022; Feng et al., 2024; Lukoff et al., 2021). Here, we define agency as “self-causality/identity” (Bennett et al., 2023), referring to the degree to which users can directly make decisions and take actions that align with their own values. Past work in HCI has explored ways to better support users’ agency, such as shaping their social media feed algorithms (Feng et al., 2024; Lam et al., 2022; Gordon et al., 2022; Stray et al., 2024) and empowering interventions against dark patterns (Lu et al., 2024). Building on this long and rich tradition, we extend the ideal of user agency to the domain of value alignment in LLM-based systems (Shapiro and Shachter, 2002; Carissa and Ghazaivi, 2023) and propose the concept of user-driven value alignment. Unlike traditional approaches (Ouyang et al., 2022; Askell et al., 2021; Bai et al., 2022; Sorensen et al., 2024), this concept highlights the active role users play in shaping AI behavior to better align with their values. This paper focuses on a case study where users identify discriminatory statements made by AI companions and actively attempt to re-align the harmful AI to reduce bias.